# Classification of unimodal isolated complete intersection singularities in positive characteristic

### v2

@@ 4. The classification of order 2 in $charF=2$
The following result is from [PPG24]:
@@
(a) $f\sim(xy,g)$ for some $f\in\mathfrak{m}^{2}$. In this case, $f\sim(xy,x^{m}+y^{n})$ for some $m,n\geq 2$, which is simple.
@@
Moreover, similar with Proposition 3.3, we can show
@@
It is easy to compute that the codimension of $\widetilde{T}_{l}(\mathcal{K}_{4}l)$ in $P_{2,4}$ is at least $2$ ($(0,xy^{3}),(0,y^{4})\notin P_{2,4}$). Therefore $\mathcal{K}-mod(f)\geq\mathcal{K}_{4}-mod(l)\geq 2$ by Proposition 2.19. ∎
@@
If $d\neq 0$, let $f_{0}=(x^{2},dy^{3})\sim(x^{2},y^{3})$ be the weighted 00-jet of $f$ with respect to $(a,d)$, where $a=(4,3),\ d=(8,9)$. Then
@@
If $d=0$ and $c\neq 0$, we still choose $f_{0}=(x^{2},cxy^{2})\sim(x^{2},xy^{2})$ be the weighted 00-jet of $f$ with respect to $(a,d)$, where $a=(4,3),\ d=(8,10)$. In this case
@@
If $g\in m^{4}$, then $f$ is equivalent to other forms in Table LABEL:table-ord2-char2.
@@
Computing the complete transversal of $j_{2}(f)=(x^{2},0)$, we have $f\sim(x^{2}+a(y)xy^{r}+b(y)y^{s},c(y)xy^{u}+d(y)y^{v})$ with $a(y),b(y),c(y),d(y)$ are units or 00. If $a(y)$ (resp. $b(y),c(y),d(y)$)$=0$, we regard as $r$ (resp. $s,u,v$) =absent=$\infty$= ∞. By Proposition 4.3, we have either $u=3$ or $v=4$.
(i) $v=4,s=3$. Then $f\sim(x^{2}+a(y)xy^{r}+b(y)y^{3},c(y)xy^{u}+d(y)y^{4})$. Take $l=(x^{2}+b_{0}y^{3},d_{0}y^{4})\sim(x^{2}+y^{3},y^{4})$ be the weighted 0-jet w.r.t. $a=(3,2),d=(6,8)$. We have $F_{a,d}^{1}\subset span\langle(0,xy^{3})\rangle+\widetilde{T}_{l}(F^{1}\mathcal{K}l)$. Therefore $f\sim(x^{2}+y^{3},y^{4}+\lambda xy^{3})$ for $\lambda\in\{0,1\}$.
(ii) $v=4,s>3$. If $r\geq 3$, then we choose $l=(x^{2},y^{4})$ be the weighted 0-jet of $f$ w.r.t. $a=(2,1),d=(4,4)$. We have $F_{a,d}^{1}\subset span\langle(xy^{3},0),(0,xy^{3})\rangle+\widetilde{T}_{l}(F^{1}\mathcal{K}l)$. Thus $f\sim(x^{2}+\lambda xy^{3},y^{4}+\mu xy^{3})$, $\lambda,\mu\in\{0,1\}$ after a scaling. That is, $f\sim k^{2}_{\lambda,\mu}$. If $r=2$, then the weighted 0-jet of $f$ w.r.t. $a=(2,1),d=(4,4)$ becomes $l=(x^{2}+a_{0}xy^{2},y^{4})\sim(x^{2}+xy^{2},y^{4})$. Computation shows that $F_{a,d}^{1}\subset\widetilde{T}_{l}(F^{1}\mathcal{K}l)$, then we have $f\sim(x^{2}+xy^{2},y^{4})\sim l^{2}$.
(iii) $v>4$. Then we have $s=3$ and then $f\sim(x^{2}+a(y)xy^{r}+b(y)y^{s},xy^{3}+d(y)y^{v})$. Choose $l=(x^{2},xy^{2})$ be the weighted 0-jet of $f$ w.r.t. $a=(1,1),d=(2,4)$. We have $F_{a,d}^{1}\subset span\langle(xy^{2},0),(y^{k},0)|\ k\geq 3\rangle+\widetilde{T}_{l}(F^{1}\mathcal{K}l)$. Then $f\sim(x^{2}+\mu xy^{2}+e(y)y^{s},xy^{3})$, $\mu\in\{0,1\},s\geq 3$. After a scaling we can furthermore assume $e(0)=1$.
@@
$\sim(x^{2}+e_{1}x^{2}y+e_{2}x^{2}y^{2}+y^{s},xy^{3})$
We apply
@@ 5. The classification of order 3
[…]
5.1. The classification of 3-jet
First choose a suitable coordinate system such that $j_{3}(f_{1})=ax^{3}+bx^{2}y+cxy^{2}+dy^{3}$ with $a,b,c,d\in F$ and $a\neq 0.$ Then $j_{3}(f_{1})\sim x^{3}+\frac{b}{a}x^{2}y+\frac{c}{a}xy^{2}+\frac{d}{a}y^{3}\sim(x-e_{1}y)(x-e_{2}y)(x-e_{3}y)\sim l_{1}l_{2}l_{3}$ since $F$ is algebraically closed, where $l_{i},\ i=1,2,3$ are linear forms in $R$.
@@ 5.1. The classification of 3-jet
I.4 If $a=0,\ b,c\neq 0,\ j_{3}(f)\sim(x^{3},x^{2}y+b^{\prime}xy^{2}).$ Let $\widetilde{x}=2b^{\prime}x,\ \widetilde{y}=y$, then $j_{3}(f)\sim(\widetilde{x}^{3},2\widetilde{x}^{2}\widetilde{y}+\widetilde{x}\widetilde{y}^{2})\sim(x^{3},x^{3}+2x^{2}y+xy^{2})\sim(x^{3},x(x+y)^{2}).$ Let $\phi(x)=x,\ \phi(y)=y-x,$ then $j_{3}(f)\sim(x^{3},xy^{2}).$
@@ I.5 If $a\neq 0,\ b^{2}\neq 3ac$, we have
$\sim(x^{3},(y+\frac{b^{\prime}}{3}x)^{3}+(c^{\prime}-\frac{b^{\prime 2}}{3})x^{2}(y+\frac{b^{\prime}}{3}x),$
@@ II.6 If $c\neq 0,\ a\neq 0,$ we have
(5.3) $j_{3}(f)$ $\sim(x^{2}(y-\frac{a^{\prime}}{3}x),y^{3}+(b^{\prime}-\frac{a^{\prime 3}}{27})x^{3})$
@@
[…]
hence such $\alpha,r,s$ exists. Then $j_{3}(f)\sim(x^{3}+xy^{2},(y+rx)(y+sx)^{2})$. Using the automorphism $\phi(x)=x+\frac{y}{s},\phi(y)=y+rx$, then it deduced to case I or II.
@@
(5.6) $\begin{split}&(x^{3},0),\ (x^{3},x^{2}y),\ (x^{3},xy^{2}),\ (x^{3},y^{3}+x^{2}y),\ (x^{3},y^{3}),\\ &(x^{2}y,0),\ (x^{2}y,xy^{2}),\ (x^{2}y,x^{3}+xy^{2}),\ (x^{3}+x^{2}y,y^{3}+\lambda x^{2}y)(\lambda\neq 0),\\ &(x^{2}y,x^{3}+y^{3}),\ (x^{3}+xy^{2},0)\end{split}$
@@ 5.2. The classification of unimodal
$u=u_{0}+u_{1}y+\dots,\ v=v0+v_{1}y+\dots,$ $p\mid 2r-2s+3$
@@
If $j_{3}(f)$ is contact equivalent to one of the following form
(5.7) $(x^{3},y^{3}+x^{2}y),\ (x^{3},y^{3}),\ (x^{2}y,x^{3}+y^{3}),\ (x^{3}+x^{2}y,y^{3}+\lambda x^{2}y),$
@@
After some computation, one can show $\mathfrak{m}^{4}\subset\mathfrak{m}\cdot\widetilde{T}_{f}(\mathcal{K}f)$ when $j_{3}(f)$ has forms in 6.2. Hence by Theorem 2.6, $j_{3}(f)$ is 3-determined. Here we compute the case when $j_{3}(f)\sim(x^{3},y^{3}+x^{2}y)$ as an example.
@@
Use a computation like Proposition 5.5, a complete transversal is given by $\{(y^{l},0),(0,x^{l})\mid l\geq 4\}.$ Hence by Corollary 2.8, we have
(5.9) $f$ $\sim(x^{2}y+\sum_{l\geq 4}a_{l}y^{l},xy^{2}+\sum_{l\geq 4}b_{l}x^{l})$
@@
Using the automorphism $\phi(x)=a(y)^{\frac{1}{2}}x$ and $\phi(y)=y$, we have $f\sim(x^{2}y+y^{r},a(y)^{\frac{1}{2}}xy^{2}+a(y)^{\frac{s}{2}}b(a(y)^{\frac{1}{2}}x)x^{s})\sim(x^{2}y+y^{r},xy^{2}+e(x,y)x^{s}),$ where $e(x,y)$ is a unit. We write $e(x,y)=\sum_{i\geq 0}e_{i}(x)y^{i},$ then $f\sim(x^{2}y+y^{r},xy^{2}+(\sum_{i\geq 0}e_{i}(x)y^{i})x^{s})\sim(x^{2}y+y^{r},xy^{2}+(\sum_{i\geq 0}e_{i}(x)y^{i})x^{s}-x^{s-2}(x^{2}y+y^{r})(\sum_{k\geq 1}e_{k}(x)y^{k-1}))\sim(x^{2}y+y^{r},xy^{2}+e_{0}(x)x^{s}-\sum_{k\geq 1}x^{s-2}y^{r+k-1})$. Repeat the operation, we get $f\sim(x^{2}y+y^{r},xy^{2}+e_{0}(x)x^{s}+d(x,y))$, where $d(x,y)=y^{\frac{rs}{2}}e^{\prime}(x,y)$ or $xy^{\frac{r(s-1)}{2}}e^{\prime}(x,y)$ depending on $s$ is even or odd. The order of $d$ is $\geq\frac{r(s-1)}{2}$.
Using the automorphism $\phi(x)=x$ and $\phi(y)=e_{0}(x)^{\frac{1}{2}}y$. Then $f\sim(e_{0}(x)^{\frac{1}{2}}x^{2}y+e_{0}(x)^{\frac{r}{2}}y^{r},xy^{2}+x^{s}+d_{1}(x,y))\sim(x^{2}y+e_{0}(x)^{\frac{r-1}{2}}y^{r},xy^{2}+x^{s}+d_{1}(x,y))$. Write $e_{0}(x)^{\frac{r-1}{2}}=\sum_{i\geq 0}u_{i}x^{i},$ then $f\sim(x^{2}y+(\sum_{i\geq 0}u_{i}x^{i})y^{r},xy^{2}+x^{s}+d_{1})\sim(x^{2}y+(\sum_{i\geq 0}u_{i}x^{i})y^{r}-y^{r-2}(xy^{2}+x^{s}+d_{1})(\sum_{k\geq 1}u_{k}x^{k-1}),xy^{2}+x^{s}+d_{1})\sim(x^{2}y+u_{0}y^{r}-\sum_{k\geq 1}u_{k}y^{r-2}x^{s+k-1}-\sum_{k\geq 1}u_{k}y^{r-2}d_{1}.$ Repeat the operation, we get $f\sim(x^{2}y+u_{0}y^{r}+d_{2},xy^{2}+x^{s}+d_{1}),$ where the order of $d_{2}$ is $\geq\frac{s(r-1)}{2}+1.$
@@
$f\sim(x^{2}y+y^{r}+\widetilde{d_{2}},xy^{2}+x^{s}+\widetilde{d_{1}}).$
Specifically, $\widetilde{d_{1}}$ has order $\frac{rs}{2}$ if $s$ is even and $\frac{r(s-1)}{2}+1$ if $s$ is odd. $\widetilde{d_{2}}$ has order $\frac{rs}{2}$ if $r$ is even and $\frac{(r-1)s}{2}+1$ if $r$ is odd.
Now we exchange the position of $x,y$ so that $r\geq s$. Let $g=j_{r}(f)=(x^{2}y+y^{r},xy^{2}+x^{s})$. Using the similar computation with the proof of (3.b.ii.2.3.3.1), we can show $\mathfrak{m}^{r+1}R^{2}\subset\mathfrak{m}\widetilde{T}_{g}(\mathcal{K}g)$ if $p\nmid rs-r-s-3$. This means $g$ is $32r-$ 3-determined by Theorem 2.6. Therefore $min\{\mathrm{ord}\widetilde{d_{1}},\mathrm{ord}\widetilde{d_{2}}\}>2r-3$ and $f\sim j_{r}(f)=(x^{2}y+y^{r},xy^{2}+x^{s})$.
@@
Denote $u_{0}(y)=e(y),\ v_{1}(y)=1-\sum_{k\geq 1}e_{k}xy^{r+k-3}.$ Then $v_{1}(y)\in 1+\mathfrak{m}^{r-2}$ and $f\sim(x^{2}y+y^{r},e_{0}x^{3}+u_{1}(y)xy^{2})\sim(x^{2}y+y^{r},e_{0}v_{1}(y)^{-1}x^{3}+xy^{2}).$ Since $v_{1}(y)\in 1+\mathfrak{m}^{r-2},$ we have $v_{1}(y)^{-1}\in 1+\mathfrak{m}^{r-2}$. Denote $u_{1}(y)=e_{0}v_{1}(y)^{-1},$ then $u_{1}(y)\in e_{0}+\mathfrak{m}^{r-2}$ and $f\sim(x^{2}y+y^{r},u_{1}(y)x^{3}+xy^{2})$. Write $u_{1}(y)=e_{0}+\sum_{k\geq r-2}v_{k}y^{k}$, we have $f\sim(x^{2}y+y^{r},(e_{0}+\sum_{k\geq r-2}v_{k}y^{k})x^{3}+xy^{2})\sim(x^{2}y+y^{r},(e_{0}+\sum_{k\geq r-2}v_{k}y^{k})x^{3}+xy^{2}-x(x^{2}y+y^{r})(\sum_{k\geq r-2}v_{k}y^{k-1}))\sim(x^{2}y+y^{r},e_{0}x^{3}+xy^{2}-\sum_{k\geq r-2}v_{k}xy^{r+k-1})\sim(x^{2}y+y^{r},e_{0}x^{3}+xy^{2}(1-\sum_{k\geq r-2}v_{k}xy^{r+k-3})).$ Denote $v_{2}(y)=1-\sum_{k\geq r-2}v_{k}xy^{r+k-3},$ then $v_{2}(y)\in 1+\mathfrak{m}^{2r-5}.$ Denote $u_{2}(y)=e_{0}v_{2}(y)^{-1}\in e_{0}+\mathfrak{m}^{2r-5},$ we have $f\sim(x^{2}y+y^{r},u_{2}(y)x^{3}+xy^{2}).$ Repeat the operation, we can get a sequence of units $u_{1}(y),u_{2}(y),\dots,u_{n}(y),\dots$ with $u_{n}(y)\in e_{0}+\mathfrak{m}^{n(r-3)+1}$ and $f\sim(x^{2}y+y^{r},u_{n}(y)x^{3}+xy^{2})$. Since $r\geq 4,$ the orders of $u_{n}(y)-e_{0}$ are strictly increasing. Note that $(x^{2}y+y^{r},e_{0}x^{3}+xy^{2})$ has finite Tjurina number (and hence is finite determined), we have $f\sim(x^{2}y+y^{r},e_{0}x^{3}+xy^{2})\sim(x^{2}y+y^{r},x^{3}+xy^{2})$ by using Remark 2.7 and applying the automorphism $\phi(x)=\alpha x,\ \phi(y)=\beta y$ where $\alpha^{2}\beta=\beta^{r}$ and $e_{0}\alpha^{3}=\alpha\beta^{2}.$

### v3

@@ 4. The classification of order 2 in $charF=2$
The following result is from [PPG25]:
@@
(a) $f\sim(xy,g)$ for some $g\in\mathfrak{m}^{2}$. In this case, $f\sim(xy,x^{m}+y^{n})$ for some $m,n\geq 2$, which is simple.
@@
Moreover, similar to Proposition 3.3, we can show
@@
It is easy to compute that the codimension of $\widetilde{T}_{l}(\mathcal{K}_{4}l)$ in $P_{2,4}$ is at least 22 ($(0,xy^{3}),(0,y^{4})\notin P_{2,4}$). Therefore $\mathcal{K}-mod(f)\geq\mathcal{K}_{4}-mod(l)\geq 2$ by Proposition 2.19. ∎
@@
If $d\neq 0$, let $f_{0}=(x^{2},dy^{3})\sim(x^{2},y^{3})$ be the weighted 0-jet of $f$ with respect to $(a,d)$, where a=(4,3),d=(8,9)a=(4,3),\ d=(8,9). Then
@@
If $d=0$ and $c\neq 0$, we still choose $f_{0}=(x^{2},cxy^{2})\sim(x^{2},xy^{2})$ be the weighted 0-jet of $f$ with respect to $(a,d)$, where a=(4,3),d=(8,10)a=(4,3),\ d=(8,10). In this case
@@
If $g\in\mathfrak{m}^{4}$, then $f$ is equivalent to other forms in Table LABEL:table-ord2-char2, that is, $j_{\lambda}^{2}$, $k_{\lambda,\mu}^{2}$, $l^{2}$, $m_{s}^{2}$, $m_{s,e_{1}}^{2}$, $n_{s}^{2}$ or $\widetilde{n}^{2}_{s,e_{1}}$.
@@
Computing the complete transversal of $j_{2}(f)=(x^{2},0)$, we have $f\sim(x^{2}+a(y)xy^{r}+b(y)y^{s},c(y)xy^{u}+d(y)y^{v})$ with $a(y),b(y),c(y),d(y)$ are units or 0. If $a(y)$ (resp. b(y),c(y),d(y)b(y),c(y),d(y))=0=0, we regard as $r$ (resp. $s,u,v$) $\infty$. By Proposition 4.3, we have either $u=3$ or $v=4$.
(i) $v=4,s=3$. Then $f\sim(x^{2}+a(y)xy^{r}+b(y)y^{3},c(y)xy^{u}+d(y)y^{4})$. Take $l=(x^{2}+b_{0}y^{3},d_{0}y^{4})\sim(x^{2}+y^{3},y^{4})$ be the weighted 0-jet with respect to $a=(3,2),d=(6,8)$. We have $F_{a,d}^{1}\subset span\langle(0,xy^{3})\rangle+\widetilde{T}_{l}(F^{1}\mathcal{K}l)$. Therefore $f\sim(x^{2}+y^{3},y^{4}+\lambda xy^{3})$ for $\lambda\in\{0,1\}$, which is $j_{\lambda}^{2}$ in Table LABEL:table-ord2-char2.
(ii) $v=4,s>3$. If $r\geq 3$, then we choose $l=(x^{2},y^{4})$ be the weighted 0-jet of $f$ with respect to $a=(2,1),d=(4,4)$. We have $F_{a,d}^{1}\subset span\langle(xy^{3},0),(0,xy^{3})\rangle+\widetilde{T}_{l}(F^{1}\mathcal{K}l)$. Thus $f\sim(x^{2}+\lambda xy^{3},y^{4}+\mu xy^{3})$, $\lambda,\mu\in\{0,1\}$ after a scaling. That is, $f\sim k^{2}_{\lambda,\mu}$ in Table LABEL:table-ord2-char2. If $r=2$, then the weighted 0-jet of $f$ with respect to $a=(2,1),d=(4,4)$ becomes $l=(x^{2}+a_{0}xy^{2},y^{4})\sim(x^{2}+xy^{2},y^{4})$. Computation shows that $F_{a,d}^{1}\subset\widetilde{T}_{l}(F^{1}\mathcal{K}l)$, then we have $f\sim(x^{2}+xy^{2},y^{4})\sim l^{2}$, which is in Table LABEL:table-ord2-char2.
(iii) $v>4$. Then we have $s=3$ and then $f\sim(x^{2}+a(y)xy^{r}+b(y)y^{s},xy^{3}+d(y)y^{v})$. Choose $l=(x^{2},xy^{2})$ be the weighted 0-jet of $f$ with respect to $a=(1,1),d=(2,4)$. We have $F_{a,d}^{1}\subset span\langle(xy^{2},0),(y^{k},0)|\ k\geq 3\rangle+\widetilde{T}_{l}(F^{1}\mathcal{K}l)$. Then $f\sim(x^{2}+\mu xy^{2}+e(y)y^{s},xy^{3})$, $\mu\in\{0,1\},s\geq 3$. After a scaling we can furthermore assume $e(0)=1$.
@@
$\sim(x^{2}+e_{1}x^{2}y+e_{2}x^{2}y^{2}+y^{s},xy^{3}),$
where $e(y)^{-1}=e_{0}+e_{1}y+e_{2}y^{2}+\dots,$ and $e_{3}$-term vanishes since $e_{3}x^{2}y^{3}$ is killed by the $xy^{3}$ term in the second component. We apply
@@ 5. The classification of order 3
In this part we assume $f=(f_{1},f_{2})\in F[x,y]^{2}$ with $\mathrm{ord}(f_{1})=3,\ \mathrm{ord}(f_{2})\geq 3.$ We also assume charF=p>3F=p>3 in this part. We begin by classifying 3-jets.
5.1. The classification of 3-jets
First choose a suitable coordinate system such that $j_{3}(f_{1})=ax^{3}+bx^{2}y+cxy^{2}+dy^{3}$ with $a,b,c,d\in F$ and $a\neq 0.$ Then $j_{3}(f_{1})\sim x^{3}+\frac{b}{a}x^{2}y+\frac{c}{a}xy^{2}+\frac{d}{a}y^{3}\sim(x-e_{1}y)(x-e_{2}y)(x-e_{3}y)\sim l_{1}l_{2}l_{3}$ since $F$ is algebraically closed, where $l_{i},\ i=1,2,3,$ are linear forms in $R$.
@@ 5.1. The classification of 3-jets
I.4 If $a=0,\ b,c\neq 0,\ j_{3}(f)\sim(x^{3},x^{2}y+b^{\prime}xy^{2}),$ where $b^{\prime}=\frac{b}{c}$. Let $\widetilde{x}=2b^{\prime}x,\ \widetilde{y}=y$, then $j_{3}(f)\sim(\widetilde{x}^{3},2\widetilde{x}^{2}\widetilde{y}+\widetilde{x}\widetilde{y}^{2})\sim(x^{3},x^{3}+2x^{2}y+xy^{2})\sim(x^{3},x(x+y)^{2}).$ Let $\phi(x)=x,\ \phi(y)=y-x,$ then $j_{3}(f)\sim(x^{3},xy^{2}).$
@@ I.5 If $a\neq 0,\ b^{2}\neq 3ac$, we have
$\sim(x^{3},(y+\frac{b^{\prime}}{3}x)^{3}+(c^{\prime}-\frac{b^{\prime 2}}{3})x^{2}(y+\frac{b^{\prime}}{3}x)),$
@@ II.6 If $c\neq 0,\ a\neq 0,$ we have
$j_{3}(f)$ $\sim(x^{2}(y-\frac{a^{\prime}}{3}x),y^{3}+(b^{\prime}-\frac{a^{\prime 3}}{27})x^{3})$ (5.3)
@@
$s^{4}+(3-v)s^{2}-2us-v=0,\ 2r=\frac{v}{s}-s,\ \alpha=rs^{2},$ (5.5)
hence such $\alpha,r,s$ exist. Then $j_{3}(f)\sim(x^{3}+xy^{2},(y+rx)(y+sx)^{2})$. Using the automorphism $\phi(x)=x+\frac{y}{s},\phi(y)=y+rx$, then it is deduced to case I or II.
@@
$\begin{split}&(x^{3},0),\ (x^{3},x^{2}y),\ (x^{3},xy^{2}),\ (x^{3},y^{3}+x^{2}y),\ (x^{3},y^{3}),\\ &(x^{2}y,0),\ (x^{2}y,xy^{2}),\ (x^{2}y,x^{3}+xy^{2}),\ (x^{3}+x^{2}y,y^{3}+\lambda x^{2}y)(\lambda\neq 0),\\ &(x^{2}y,x^{3}+y^{3}),\ (x^{3}+xy^{2},0).\end{split}$} (5.6)
@@ II $(x^{3},y^{3}+x^{2}y)$
$\widetilde{P}_{r,s}$ $(x^{3}+xy^{3}+uxy^{r}+vy^{s},x^{2}y+y^{4}),\ where$ $r\geq 4,\ s\geq 5,$
$u=u_{0}+u_{1}y+\dots,\ {v=v_{0}+v_{1}y+\dots,}$ $p\mid 2r-2s+3$
@@
If $j_{3}(f)$ is contact equivalent to one of the following forms
$(x^{3},y^{3}+x^{2}y),\ (x^{3},y^{3}),\ (x^{2}y,x^{3}+y^{3}),\ (x^{3}+x^{2}y,y^{3}+\lambda x^{2}y)(\lambda\neq 0),$ (5.7)
@@
After some computation, one can show $\mathfrak{m}^{4}\subset\mathfrak{m}\cdot\widetilde{T}_{f}(\mathcal{K}f)$ when $j_{3}(f)$ has forms in (5.2). Hence by Theorem 2.6, $j_{3}(f)$ is 3-determined. Here we compute the case when $j_{3}(f)\sim(x^{3},y^{3}+x^{2}y)$ as an example.
@@
Use a computation like Proposition 5.5, a complete transversal is given by $\{(y^{l},0),(0,x^{l})\mid l\geq 4\}.$ Hence by Theorem 2.8, we have
$f$ $\sim(x^{2}y+\sum_{l\geq 4}a_{l}y^{l},xy^{2}+\sum_{l\geq 4}b_{l}x^{l})$ (5.9)
@@
Using the automorphism $\phi(x)=a(y)^{\frac{1}{2}}x$ and $\phi(y)=y$, we have
$f$ $\sim(x^{2}y+y^{r},a(y)^{\frac{1}{2}}xy^{2}+a(y)^{\frac{s}{2}}b(a(y)^{\frac{1}{2}}x)x^{s})$ (5.10)
$\sim(x^{2}y+y^{r},xy^{2}+e(x,y)x^{s}),$
where $e(x,y)$ is a unit. We write $e(x,y)=\sum_{i\geq 0}e_{i}(x)y^{i},$ then
$f$ $\sim(x^{2}y+y^{r},xy^{2}+(\sum_{i\geq 0}e_{i}(x)y^{i})x^{s})$ (5.11)
$\sim(x^{2}y+y^{r},xy^{2}+(\sum_{i\geq 0}e_{i}(x)y^{i})x^{s}-x^{s-2}(x^{2}y+y^{r})(\sum_{k\geq 1}e_{k}(x)y^{k-1}))$
$\sim(x^{2}y+y^{r},xy^{2}+e_{0}(x)x^{s}-\sum_{k\geq 1}x^{s-2}{e_{k}(x)}y^{r+k-1}).$
Note that the order of $x$ changes from $s$ to s2s-2, and the order of $y$ changes from 0 to $r$. Repeat the operation, we get $f\sim(x^{2}y+y^{r},xy^{2}+e_{0}(x)x^{s}+d(x,y))$, where $d(x,y)=y^{\frac{rs}{2}}e^{\prime}(x,y)$ or $xy^{\frac{r(s-1)}{2}}e^{\prime}(x,y)$ depending on $s$ is even or odd, and $e^{\prime}(x,y)=\sum_{k\geq 1}e_{k}(x)y^{k-1}$, which is a unit in $F[[x,y]]$. The order of $d$ is $\geq\frac{r(s-1)}{2}+1$.
Using the automorphism $\phi(x)=x$ and $\phi(y)=e_{0}(x)^{\frac{1}{2}}y$. Then
$f$ $\sim(e_{0}(x)^{\frac{1}{2}}x^{2}y+e_{0}(x)^{\frac{r}{2}}y^{r},xy^{2}+x^{s}+d_{1}(x,y))$ (5.12)
$\sim(x^{2}y+e_{0}(x)^{\frac{r-1}{2}}y^{r},xy^{2}+x^{s}+d_{1}(x,y)),$
where $d_{1}(x,y)=d(x,e_{0}(x)^{\frac{1}{2}}y)$. Write $e_{0}(x)^{\frac{r-1}{2}}=\sum_{i\geq 0}u_{i}x^{i},$ then
$f$ $\sim(x^{2}y+(\sum_{i\geq 0}u_{i}x^{i})y^{r},xy^{2}+x^{s}+d_{1}(x,y))$ (5.13)
$\sim(x^{2}y+(\sum_{i\geq 0}u_{i}x^{i})y^{r}-y^{r-2}(xy^{2}+x^{s}+d_{1}(x,y))(\sum_{k\geq 1}u_{k}x^{k-1}),xy^{2}+x^{s}+d_{1}(x,y))$
$\sim(x^{2}y+u_{0}y^{r}-\sum_{k\geq 1}u_{k}y^{r-2}x^{s+k-1}-\sum_{k\geq 1}u_{k}y^{r-2}d_{1}(x,y),xy^{2}+x^{s}+d_{1}(x,y)).$
Repeat the operation, we get $f\sim(x^{2}y+u_{0}y^{r}+d_{2}(x,y),xy^{2}+x^{s}+d_{1}(x,y)),$ where the order of $d_{2}(x,y)$ is $\geq\frac{s(r-1)}{2}+1.$
@@
$f\sim(x^{2}y+y^{r}+\widetilde{d_{2}}(x,y),xy^{2}+x^{s}+\widetilde{d_{1}}(x,y)).$
Specifically, $\widetilde{d_{1}}(x,y)$ has order $\frac{rs}{2}$ if $s$ is even and $\frac{r(s-1)}{2}+1$ if $s$ is odd. $\widetilde{d_{2}}(x,y)$ has order $\frac{rs}{2}$ if $r$ is even and $\frac{(r-1)s}{2}+1$ if $r$ is odd.
Now we exchange the position of $x,y$ so that $r\geq s$. Let $g=j_{r}(f)=(x^{2}y+y^{r},xy^{2}+x^{s})$. Using the similar computation as the proof of (3.b.ii.2.3.3.1) in Proposition 3.6, we can show $\mathfrak{m}^{r+1}R^{2}\subset\mathfrak{m}\widetilde{T}_{g}(\mathcal{K}g)$. This means $g$ is (2r3)(2r-3)-determined by Theorem 2.6. Since $s\geq 4$, $min\{\mathrm{ord}\widetilde{d_{1}}(x,y),\mathrm{ord}\widetilde{d_{2}}(x,y)\}>2r-3$. Therefore $f\sim j_{2r-3}(f)=(x^{2}y+y^{r},xy^{2}+x^{s})$.
@@
Denote $u_{0}(y)=e(y),\ v_{1}(y)=1-\sum_{k\geq 1}e_{k}xy^{r+k-3}.$ Then $v_{1}(y)\in 1+\mathfrak{m}^{r-2}$ and
$f$ $\sim(x^{2}y+y^{r},e_{0}x^{3}+u_{1}(y)xy^{2})$ (5.15)
$\sim(x^{2}y+y^{r},e_{0}v_{1}(y)^{-1}x^{3}+xy^{2}).$
Since $v_{1}(y)\in 1+\mathfrak{m}^{r-2},$ we have $v_{1}(y)^{-1}\in 1+\mathfrak{m}^{r-2}$. Denote $u_{1}(y)=e_{0}v_{1}(y)^{-1},$ then $u_{1}(y)\in e_{0}+\mathfrak{m}^{r-2}$ and $f\sim(x^{2}y+y^{r},u_{1}(y)x^{3}+xy^{2})$. Write
$u_{1}(y)=e_{0}+\sum_{k\geq r-2}v_{k}y^{k},$
we have
$f$ $\sim(x^{2}y+y^{r},(e_{0}+\sum_{k\geq r-2}v_{k}y^{k})x^{3}+xy^{2})$ (5.16)
$\sim(x^{2}y+y^{r},(e_{0}+\sum_{k\geq r-2}v_{k}y^{k})x^{3}+xy^{2}-x(x^{2}y+y^{r})(\sum_{k\geq r-2}v_{k}y^{k-1}))$
$\sim(x^{2}y+y^{r},e_{0}x^{3}+xy^{2}-\sum_{k\geq r-2}v_{k}xy^{r+k-1})$
$\sim(x^{2}y+y^{r},e_{0}x^{3}+xy^{2}(1-\sum_{k\geq r-2}v_{k}y^{r+k-3})).$
Denote $v_{2}(y)=1-\sum_{k\geq r-2}v_{k}y^{r+k-3},$ then $v_{2}(y)\in 1+\mathfrak{m}^{2r-5}.$ Denote $u_{2}(y)=e_{0}v_{2}(y)^{-1}\in e_{0}+\mathfrak{m}^{2r-5},$ we have $f\sim(x^{2}y+y^{r},u_{2}(y)x^{3}+xy^{2}).$
Repeat the operation, we can get a sequence of units $u_{1}(y),u_{2}(y),\dots,u_{n}(y),\dots$ with $u_{n}(y)\in e_{0}+\mathfrak{m}^{n(r-3)+1}$ and $f\sim(x^{2}y+y^{r},u_{n}(y)x^{3}+xy^{2})$.
Since $r\geq 4,$ the orders of $u_{n}(y)-e_{0}$ are strictly increasing. Note that $(x^{2}y+y^{r},e_{0}x^{3}+xy^{2})$ has finite Tjurina number (and hence is finite determined), we have $f\sim(x^{2}y+y^{r},e_{0}x^{3}+xy^{2})\sim(x^{2}y+y^{r},x^{3}+xy^{2})$ by using Remark 2.7 and applying the automorphism $\phi(x)=\alpha x,\ \phi(y)=\beta y$ where $\alpha^{2}\beta=\beta^{r}$ and $e_{0}\alpha^{3}=\alpha\beta^{2}.$